OpenStudy (anonymous):

\[f(x)=8x ^{2}+16x-3\] vertex?

myininaya (myininaya):

put it in vertex form f(x)=a(x-h)^2+k

myininaya (myininaya):

f(x)=8(x^2+2x)-3 f(x)=8(x^2+2x+1)-3-8 f(x)=8(x+1)^2-11

myininaya (myininaya):

is that what you got when you tried it?